# Contractive kinetic Langevin samplers beyond global Lipschitz continuity

### v1

@@ 6.4 Proof of Log-Sobolev inequality for the OBABO scheme
a.s which implies the $W_{\infty}$ contraction of the Markov operator $P_{a,b}$. By Theorem 2.2 in Kuwada (2010) one obtains the result.
@@
See proof of Theorem 23, point 1, in Monmarché (2021). ∎
@@ 6.5 Proof of Log-Sobolev inequality for the exponential Euler scheme
$\nabla\bar{F}(\varphi,\psi)=\begin{pmatrix}\partial_{\varphi}\bar{F}_{\varphi}&\partial_{\varphi}\bar{F}_{\psi}\\ \partial_{\psi}\bar{F}_{\varphi}&\partial_{\psi}\bar{F}_{\psi}\end{pmatrix}=\begin{pmatrix}\frac{1+e^{-\gamma\lambda}}{2}I_{d}-\frac{\lambda-\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)&\frac{1-e^{-\gamma\lambda}}{2}I_{d}-\frac{\lambda+\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)\\ \frac{1-e^{-\gamma\lambda}}{2}I_{d}&\frac{1+e^{-\gamma\lambda}}{2}I_{d}\end{pmatrix}$.
@@
$\nabla\bar{F}(\varphi,\psi)=A+B=\frac{1}{2}\begin{pmatrix}(1+\eta)I_{d}&(1-\eta)I_{d}-b\nabla h_{\lambda}(\varphi)\\ (1-\eta)I_{d}&(1+\eta)I_{d}\par\end{pmatrix}+\begin{pmatrix}\frac{\lambda-\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)&0_{d}\\ 0_{d}&0_{d}\end{pmatrix}$
@@
Proof of (1).
Noticing that $t(x)=1$ and $s(x)=0\quad\forall x\in B(0,r_{\lambda}-2),$ it follows that
$h_{\lambda}(x)=g_{\lambda}(x)+mx=g(x)+mx=h(x)\quad\forall x\in B(0,r_{\lambda}-2).$
Proof of (3).
We define
$A_{1}=\bar{B}(0,r_{\lambda}-2)\quad A_{2}=\bar{B}(0,r_{\lambda}-1)/{B}(0,r_{\lambda}-2)\quad A_{3}=\bar{B}(0,r_{\lambda})/B_{r_{\lambda}-1},\quad A_{4}=B(0,r_{\lambda})^{c}$
On $A_{1}$ there holds since $g_{\lambda}=g$
$|g_{\lambda}(x)-g_{\lambda}(y)|\leq Lip_{B_{r}}|x-y|.$
For $x,y\in A_{2}$ $Lip(s(x))=1$ and $t(x)=1$ on $A_{2}$ one obtains
For $x,y\in A_{3}$ since $Lip(t)=1$ and
$|g(x)|\leq R_{\lambda}\quad\forall x\in A_{3}$
then,
$\geq R_{\lambda}^{2}(|x|-|y|)^{2}$
$\geq 0$
so for each set $A_{i}$ there holds
$|g_{\lambda}(x)-g_{\lambda}(y)|$ $=|g_{\lambda}(l_{0})-g_{\lambda}(l_{k})|$
$\leq\sum_{i=1}^{k}|g(l_{i-1})-g(l_{i})|$
$\leq(Lip_{B_{r}}+R_{\lambda}r_{\lambda})sum_{i=1}^{k}|l_{i}-l_{i-1}|$
$=(Lip_{B_{r}}+R_{\lambda}r_{\lambda})|x-y|$
where the last line holds since the segments are collinear. ∎
Proof of (4).
$\geq R_{\lambda}(s(x)|x|-s(y)|y|)(|x|-|y|)$
$\geq 0.$
For $x,y\in A_{3},$
$\langle g_{\lambda}(x)-g_{\lambda}(y),x-y\rangle$ $=R_{\lambda}|x-y|^{2}+t(y)\langle g(x)-g(y),x-y\rangle+(t(x)-t(y))\rangle g(x),x-y\rangle$
[…]
$\geq(R-|g(x)|)|x-y|^{2}\geq 0$
For $x,y\in A_{4}$,
$\langle g_{\lambda}(x)-g_{\lambda}(y),x-y\rangle=\sum_{i=1}^{k}\langle g_{\lambda}(l_{i})-g_{\lambda}(l{i-1}),l_{i}-l_{i-1}\rangle\geq 0.$
@@
$\tilde{x}_{\lambda}$ $=\tilde{x}_{0}+\int_{0}^{\lambda}\psi_{0}(s)ds\tilde{Y_{0}}-h_{\lambda}(\tilde{x}_{0})\int_{0}^{\lambda}\int_{0}^{s}e^{-\gamma(s-u)}duds+\sqrt{2\gamma}\int_{0}^{\lambda}e^{-\gamma t}\int_{0}^{t}e^{\gamma s}dW_{s}dt$ (19)
@@
$\mathbb{E}\left(e^{-\lambda\gamma}\int_{0}^{\lambda}e^{\gamma s}dWs\right)\left(\frac{1}{\gamma}\int_{0}(1-e^{\gamma(s-\lambda)})dW_{s}\right)^{T}$ $=\frac{1}{\gamma}\left(\int_{0}^{\lambda}e^{\gamma(s-\lambda)}(1-e^{\gamma(s-\lambda)})ds\right)I_{d}$
$=\int_{0}^{\lambda}\psi_{0}(s)\psi_{1}(s)ds\quad I_{d}$
We have essentially proved that the terms $\int_{0}^{\lambda}e^{-\gamma t}M_{t}dt$ appearing in (19) and $e^{-\lambda\gamma}\int_{0}^{\lambda}e^{\gamma s}dW_{s}$ in (18) follow Gaussian distributions with the required cross-covariance matrix.
@@ 7 Convergence of the Monotonic polygonal Stochastic exponential scheme
$\mathbb{E}|\tilde{x}^{X,n,\lambda}_{t}-p_{t}|^{2}\leq\hat{C}\lambda^{5}+\frac{1}{\lambda}\int_{n\lambda}^{t}\int_{n\lambda}^{s}\mathbb{E}|\tilde{x}^{X,n,\lambda}_{u}-p_{u}|^{2}duds\quad\forall t\in[n\lambda,(n+1)\lambda]$
@@ 8.1 Comparing the BAB steps to a Hamiltonian system
$\hat{x}_{t}^{\prime}=\hat{v}_{t},\quad\hat{v}_{t}^{\prime}=-\nabla\left(\hat{x}_{t}\right),\quad\hat{x}_{0}=x,\quad\hat{v}_{0}=v,$ (34)
@@
$\sup_{t\in[0,\lambda]}|\hat{x}_{t}|^{p}+|\hat{v}_{t}|^{p}\leq C(1+|u(x)|^{p}+|v|^{p})\quad\forall p\in\mathbb{N}.$
@@
$\mathbb{E}|\hat{x}_{t}|^{p}+\mathbb{E}|\hat{v}_{t}|^{p}\leq\hat{C}_{p}\quad\forall t\in[0,\lambda]$
@@
$x_{t}^{\prime}=w_{t},\quad w_{t}^{\prime}=-h_{\lambda}(x),\quad x_{0}=x,\quad w_{0}=v$
@@
$x_{t}=x+tv-\frac{t^{2}}{2}h_{\lambda}(x),\quad w_{t}=v-th_{\lambda}(x),$
@@ Appendix A Appendix
Lemma 19 ( Raginsky et al. (2017), Lemma 3, adapted).
@@
Notice that in the proof of Lemma 3 in Raginsky et al. (2017) only the dissipativity condition is used which is a trivial consequence of strong convexity. ∎
@@
[Hargé (2004), Theorem 1.1] Let $X$ follow $\mathcal{N}(m_{0},\Sigma)$ with density $\varphi$, and let $Y$ have density $\varphi f$ where $f$ is a log-concave function. Then for any convex map $g$ there holds:

### v2

@@ 6.4 Proof of Log-Sobolev inequality for the OBABO scheme
a.s which implies the $W_{\infty}$ contraction of the Markov operator $P_{a,b}$. By Theorem 2.2 in kuwada2010duality one obtains the result.
@@
See proof of Theorem 23, point 1, in monmarche2021high. ∎
@@ 6.5 Proof of Log-Sobolev inequality for the exponential Euler scheme
$\nabla\bar{F}(\varphi,\psi)=\begin{pmatrix}\partial_{\varphi}\bar{F}_{\varphi}&\partial_{\varphi}\bar{F}_{\psi}\\ \partial_{\psi}\bar{F}_{\varphi}&\partial_{\psi}\bar{F}_{\psi}\end{pmatrix}\\ =\begin{pmatrix}\frac{1+e^{-\gamma\lambda}}{2}I_{d}-\frac{\lambda-\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)&\frac{1-e^{-\gamma\lambda}}{2}I_{d}-\frac{\lambda+\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)\\ \frac{1-e^{-\gamma\lambda}}{2}I_{d}&\frac{1+e^{-\gamma\lambda}}{2}I_{d}\end{pmatrix}$.
@@
$\nabla\bar{F}(\varphi,\psi)=A+B\\ =\frac{1}{2}\begin{pmatrix}(1+\eta)I_{d}&(1-\eta)I_{d}-b\nabla h_{\lambda}(\varphi)\\ (1-\eta)I_{d}&(1+\eta)I_{d}\end{pmatrix}+\begin{pmatrix}\frac{\lambda-\gamma^{-1}(1-e^{-\gamma\lambda})}{\gamma}\nabla h_{\lambda}(\varphi)&0_{d}\\ 0_{d}&0_{d}\end{pmatrix}$
@@
$\tilde{x}_{\lambda}$ $=\tilde{x}_{0}+\int_{0}^{\lambda}\psi_{0}(s)ds\tilde{Y_{0}}-h_{\lambda}(\tilde{x}_{0})\int_{0}^{\lambda}\int_{0}^{s}e^{-\gamma(s-u)}duds$ (20)
$\hskip 18.49988pt+\sqrt{2\gamma}\int_{0}^{\lambda}e^{-\gamma t}\int_{0}^{t}e^{\gamma s}dW_{s}dt$
@@
$\mathbb{E}\left(e^{-\lambda\gamma}\int_{0}^{\lambda}e^{\gamma s}dWs\right)\left(\frac{1}{\gamma}\int_{0}(1-e^{\gamma(s-\lambda)})dW_{s}\right)^{T}\\ =\frac{1}{\gamma}\left(\int_{0}^{\lambda}e^{\gamma(s-\lambda)}(1-e^{\gamma(s-\lambda)})ds\right)I_{d}=\int_{0}^{\lambda}\psi_{0}(s)\psi_{1}(s)ds\qquad I_{d}$
We have proved that the terms $\int_{0}^{\lambda}e^{-\gamma t}M_{t}dt$ appearing in (20) and $e^{-\lambda\gamma}\int_{0}^{\lambda}e^{\gamma s}dW_{s}$ in (19) follow Gaussian distributions with the required cross-covariance matrix.
@@ 7 Convergence of the Monotonic polygonal Stochastic exponential scheme
$\mathbb{E}|\tilde{x}^{X,n,\lambda}_{t}-p_{t}|^{2}\leq\hat{C}\lambda^{5}+\frac{1}{\lambda}\int_{n\lambda}^{t}\int_{n\lambda}^{s}\mathbb{E}|\tilde{x}^{X,n,\lambda}_{u}-p_{u}|^{2}duds\qquad\forall t\in[n\lambda,(n+1)\lambda]$
@@ 8.1 Comparing the BAB steps to a Hamiltonian system
$\hat{x}_{t}^{\prime}=\hat{v}_{t},\qquad\hat{v}_{t}^{\prime}=-\nabla\left(\hat{x}_{t}\right),\qquad\hat{x}_{0}=x,\qquad\hat{v}_{0}=v,$ (35)
@@
$\sup_{t\in[0,\lambda]}|\hat{x}_{t}|^{p}+|\hat{v}_{t}|^{p}\leq C(1+|u(x)|^{p}+|v|^{p})\qquad\forall p\in\mathbb{N}.$
@@
$\mathbb{E}|\hat{x}_{t}|^{p}+\mathbb{E}|\hat{v}_{t}|^{p}\leq\hat{C}_{p}\qquad\forall t\in[0,\lambda]$
@@
$x_{t}^{\prime}=w_{t},\qquad w_{t}^{\prime}=-h_{\lambda}(x),\qquad x_{0}=x,\qquad w_{0}=v$
@@
$x_{t}=x+tv-\frac{t^{2}}{2}h_{\lambda}(x),\qquad w_{t}=v-th_{\lambda}(x),$
@@ Appendix A Appendix
Lemma 19 ( raginsky, Lemma 3, adapted).
@@
Notice that in the proof of Lemma 3 in raginsky only the dissipativity condition is used which is a trivial consequence of strong convexity. ∎
@@
[harge2004convex, Theorem 1.1] Let $X$ follow $\mathcal{N}(m_{0},\Sigma)$ with density $\varphi$, and let $Y$ have density $\varphi f$ where $f$ is a log-concave function. Then for any convex map $g$ there holds: